# Rational powers, invariant ideals, and the summation formula

### v2

@@ Abstract.
We study rational powers and Rees valuations of invariant ideals, providing explicit descriptions in terms of polyhedra and lattice points. This allows us to show that a version of Mustaţă-Takagi’s summation formula for rational powers holds for these ideals. Moreover, for arbitrary ideals over the complex numbers, we prove a weaker version of this formula that holds for sufficiently large rational numbers.
@@ 1. Introduction
Rational powers of ideals and their relationship with Rees valuations play a pivotal role in the proofs of various results in algebraic geometry and commutative algebra [33, 36, 17, 12, 11, 43, 24, 40, 3, 23]. Originally considered by Samuel and Rees in the 1950s [53, 48], the notion of rational powers $\overline{I^{w}}$ of an ideal $I$ in a Noetherian domain $R$ was introduced by Lejeune-Jalabert and Teissier [42, §4.2]. Defined as $\overline{I^{w}}:=\cap_{v}\{x\mid v(x)\geqslant wv(I)\}$, where the intersection is taken over the discrete valuations of $\operatorname{Quot}(R)$ that are nonnegative in $R$, this construction provides a powerful tool for understanding the structure of filtrations of ideals, especially integral closures, symbolic powers, and multiplier ideals. Despite its historical roots, rational powers have seen limited development, as noted by Knutson, who observed that this “beautiful construction has sat essentially unused in the literature for over 50 years now" [39].
In the definition of rational powers one only needs to consider the intersection over finitely many valuations, the Rees valuations. The latter are defined as the order valuations corresponding to prime divisors of exceptional divisors of blowups. Therefore, knowing the Rees valuations of an ideal provides key insights on the structure of its blowup and a complete description of its rational powers; see [48], [41, Example 9.6.3], [35, Chapter 10]. However, these valuations are only known in limited cases. For example, for monomial ideals where they correspond to supporting hyperplanes of Newton polyhedra [35, Proposition 1.4.6, Theorem 10.3.5].
As one of the primary contributions of this paper, we introduce the concept of Rees package, which encapsulates the description of both Rees valuations and rational powers in terms of polyhedral data (see Section missing 3). We show that several classes of invariant ideals of great interest in the literature, including monomial ideals in affine semigroup rings, and sums of products of determinantal ideals of generic, symmetric, and skew-symmetric matrices, and products of determinantal ideals of Hankel matrices, admit such a Rees package.
@@
The rational powers and Rees valuations of the following classes of ideals can be explicitly described in terms of lattice points and supporting hyperplanes of polyhedra.
@@
    (Theorem 3.2, §missing 3.2) Sums of products of determinantal ideals of a generic matrix $X_{m\times n}$ and $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[X_{m\times n}]$.
@@
    (Theorem 3.3, §missing 3.3) Sums of products of determinantal ideals of a generic symmetric matrix $Y_{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[Y_{n\times n}]$.
@@
    (Theorem 3.4, §missing 3.4) Sums of products of ideals of Pfaffians of a skew-symmetric matrix $Z_{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[Z_{n\times n}]$.
@@
    (Theorem 3.5) Products of determinantal ideals of a Hankel matrix of variables.
Theorem A is obtained via the representation-theoretic methods developed by DeConcini-Eisenbud-Procesi [15] in the generic case and later extended to the other cases [2, 1, 13].
Multiplier ideals of ideal sheaves have been of great interest for decades due to their relation with vanishing theorems and the information they encode about singularities of higher dimensional algebraic varieties [41, Chapter 9]. They have also been applied to obtain deep results in commutative algebra, such as Ein-Lazarsfeld-Smith’s relation between the powers and symbolic powers of an ideal [25], and Lipman’s Briançon–Skoda Theorem [44]. These results came as consequences of strong connections between multiplier ideals, symbolic powers, and integral closures. In [47], Mustaţă showed a summation formula for multiplier ideals of pairs of nonzero ideal sheaves, and this formula was later generalized by Takagi [55] and by Jow-Miller [38]. Mustaţă-Takagi summation formula raises the natural question of whether similar identities hold for symbolic powers and rational powers of ideals. The former was established in [30, 29] and the latter in [4] for monomial ideals. Thus we propose the following conjecture (cf. [4, Question 2.4]).
@@
From Theorem A it follows that for invariant ideals, the Rees valuations of $IT+JT$ are obtained by pairing those of $I$ and $J$ (see Section missing 4). In Section missing 4 we ask weather this is always the case over algebraically closed fields. Theorem A also allows us to settle Section missing 1 for invariant ideals without restrictions on the field or on $w$.
Theorem B (Theorem 4.1, Section missing 4).
Adopt the notation in Section missing 1 with $\mathbb{k}$ an arbitrary field. The conclusion of Section missing 1 holds if $I$ and $J$ belong to any of the classes of ideals in Theorem A, and not necessarily both from the same class.
In our next theorem, we prove an inclusion that can be seen as a weaker version of Section missing 1 for arbitrary ideals and $w\gg 0$, over the complex numbers (see Section missing 5).
Theorem C (Theorem 5.1).
Following the notation in Section missing 1 with $\mathbb{k}=\mathbb{C}$, there exists a $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
The proof of Theorem C is influenced by that of [47, Theorem 0.3]. The first step is to define rational powers of ideal sheaves in a normal complex varieties inspired by [42, Définition 4.2.2] and [41, Definition 9.6.2]. Next, we prove the statement for ideals in discrete valuation rings. Finally, with a careful use of Serre’s vanishing theorem we obtain the desired statement.
@@ 2. Rational powers and Rees valuations
In this section we include relevant facts (old and new) about integral closures, rational powers, and Rees valuations. Some of the main references in this subject are [35, 57, 51]. Although these constructions and their properties hold generally, we focus our exposition to the following setup where all of our results occur.
@@
If $w$ is an integer, the statement follows from [35, Proposition 10.5.5]. Now let $w$ be arbitrary and write $w=\frac{p}{q}$. If $x\in\overline{I^{w}}$, then $(W^{-1}x)^{q}=W^{-1}x^{q}\in W^{-1}\overline{I^{p}}=\overline{(W^{-1}I)^{p}}$ and so $W^{-1}x\in\overline{(W^{-1}I)^{w}}$. On the other hand, if $y\in\overline{(W^{-1}I)^{w}}$, then $y^{q}\in\overline{(W^{-1}I)^{p}}=W^{-1}\overline{I^{p}}$. Therefore there exists $a\in W$ such that $ay^{q}\in\overline{I^{p}}$. The latter implies $(ay)^{q}\in\overline{I^{p}}$ and so $ay\in\overline{I^{w}}$. Thus $y\in W^{-1}\overline{I^{w}}$, finishing the proof. ∎
A valuation is discrete if it is integer valued. Integral closures and rational powers can also be defined via discrete valuations: by [35, Theorem 6.8.3(3)] we have that $x\in\overline{I}$ if and only if $x\in IV$ for every discrete valuation ring $R\subseteq V\subseteq\operatorname{Quot}(R)$. Moreover, the latter condition can be verified by using finitely many valuations, the Rees valuations of $I$.
@@
Let ${\operatorname{\mathscr{RV}}}(I)=\{v_{1},\dots,v_{r}\}$ and set $e:={\operatorname{lcm}}(v_{1}(I),\dots,v_{r}(I))$ (or any multiple of it). For every $w\in{\mathbb{Q}}_{\geqslant 0}$ and $n:=\lceil we\rceil$ we have $\overline{I^{w}}=\overline{I^{\frac{n}{e}}}$ (see [35, Proposition 10.5.5]). Moreover, $\overline{I^{\frac{1}{e}}}=\sqrt{I}$ [35, Example 10.5.4].
@@
The second equality is proved in [35, Discussion 10.1.3] and the inclusion $\supseteq$ in the first one is trivial. It remains to show the first inclusion $\subseteq$. For each $V\in{\operatorname{\mathscr{RV}}}(I)$, let $\eta_{V}$ be its maximal ideal. From Section missing 2 we obtain $\overline{I^{w}}=\cap_{V\in{\operatorname{\mathscr{RV}}}(I)}(\eta_{V}^{\lceil wv(I)\rceil}\cap R)$. The result follows as each ideal $\eta_{V}^{\lceil wv(I)\rceil}\cap R$ is $\mathfrak{p}_{V}$-primary. ∎
@@
The implication (i) $\Rightarrow$ (ii) follows from Theorem 2.1, and (ii) $\Rightarrow$ (iii) from Section missing 2. It remains to show (iii) $\Rightarrow$ (i), but this is content of [35, Theorem 10.1.6], finishing the proof. ∎
@@
For a particular example, consider $R$ a polynomial ring over a field and $H_{1},\ldots,H_{t}$ subspaces of $R_{1}$. For each $i$ let $I_{i}$ be the ideal generated by $H_{i}$ and set $I=I_{1}\cdots I_{t}$. The powers of $I$ are all integrally closed and irredundant primary decompositions of them having the same format above can be found in [14, Theorem 3.2]. Therefore, these primary decompositions can be used to explicitly describe the Rees valuations and rational powers of $I$.
@@
We begin with (1). By [35, Theorem 2.3.2], the algebra $C$ is ${\mathbb{Z}}_{\geqslant 0}$-graded and its $e$th-Veronese is the integral closure of $A_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{n}u^{en}$ in $B_{e}:=R[u^{e}]$, which is $C_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{n}}u^{en}$ by [35, Proposition 5.2.1]. Write $C=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}J_{n}u^{n}$ and fix $n$. If $x\in J_{n}$, then $x^{e}\in J_{en}=\overline{I^{n}}$, and so $x\in\overline{I^{\frac{n}{e}}}$. On the other hand, if $x\in\overline{I^{\frac{n}{e}}}$, then $x^{e}\in\overline{I^{n}}$, and so $(xu^{n})^{e}\in C$. Thus, $xu^{n}$ is integral over $A$ and then $xu^{n}\in C$. It follows that $x\in J_{n}$.
We continue with (2). Note that $A=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{\lceil\frac{n}{e}\rceil}u^{n}$ and is Noetherian. By Rees’ theorem [49] (see also [35, Theorem 9.1.2]), there exists $k\in{\mathbb{Z}}_{>0}$ such that, for every integer $n\geqslant ek$, we have
@@
Finally we prove (3). By (2) and [26, Lemma 13.10], there exists $m\in{\mathbb{Z}}_{>0}$ such that $\overline{I^{mt}}=\left(\overline{I^{m}}\right)^{t}$ for every $t\in{\mathbb{Z}}_{\geqslant 0}$. Set $C_{m}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{mt}}u^{emt}$. From (2), it follows that $C$ is a finitely generated $C_{m}$-algebra, and since $C_{m}\subset C$ is an integral extension, by [35, Lemma 2.1.9], $C$ is a finitely generated $C_{m}$-module. The statement now follows by considering $r\in{\mathbb{Z}}_{>0}$ to be the largest degree of an element in a finite set of generators of $C$ as a $C_{m}$-module. ∎
We continue with the following statement that allows for base change in rational powers (cf. [35, Corollary 19.5.2]). A ring homomorphism $f:R\to S$ is normal if it is flat and if for every $\mathfrak{p}\in{\operatorname{Spec\,}}(R)$ and field extension $\mathbb{K}$ of $R_{\mathfrak{p}}/\mathfrak{p}R_{\mathfrak{p}}$ the ring $\mathbb{K}\otimes_{R}S$ is normal.
@@
Let $e_{1}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(I)\}\right)$, $e_{2}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(IS)\}\right)$, and $e:={\operatorname{lcm}}(e_{1},e_{2})$. Let $u,A,B,C$ be as in Section missing 2 defined using the latter $e$. By [35, Theorem 19.5.1], $C\otimes_{R}S$ is the integral closure of $A\otimes_{R}S$ in $B\otimes_{R}S$. Under the isomorphism $B\otimes_{R}S\cong S[u]$, $A\otimes_{R}S$ corresponds to $S[(IS)u,(IS)u^{2},\ldots,(IS)u^{e}]$ and $C\otimes_{R}S$ to $\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}(\overline{I^{\frac{n}{e}}}S)u^{n}$. The result follows by applying Section missing 2(1) to $IS$. ∎
@@
In this subsection we introduce the new notion of rational powers of ideal sheaves in normal complex varieties inspired by [42, Définition 4.2.2] and [41, Definition 9.6.2].
@@
Let $X$ be a normal complex variety (integral, separated, and of finite type) and let $\mathcal{I}$ be a nonzero $\mathcal{O}_{X}$-ideal.
@@
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor. Note that $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. The integral closure of $\mathcal{I}$, denoted by $\overline{\mathcal{I}}$, is the $\mathcal{O}_{X}$-ideal ¯$\overline{\mathcal{I}}:=v_{*}\left(\mathcal{O}_{X^{+}}(-E)\right)$ [41, Definition 9.6.2]. Moreover, for $w\in\mathbb{Q}_{\geqslant 0}$, we call the $\mathcal{O}_{X}$-ideal
@@
If $X$ in Section missing 2 is affine, so that $\mathcal{I}=\tilde{I}$ for some $\mathbb{C}[X]$-ideal $I$, then from Section missing 2 it follows that $\overline{I^{w}}=\Gamma(X,\overline{\mathcal{I}^{w}})$ are the rational powers of $I$ as in Section missing 2.
@@
We have the following alternative way of computing rational powers similar to [41, Remark 9.6.4] and [42, Théorème 2.1]. We use the definition of projective morphism from [22], which is more general than the one of [31] (see [54, tag 01W7, Lemma 29.43.3]), and includes blowups of arbitrary varieties along ideal sheaves [54, tag 01OF, Lemma 31.32.13].
@@
By the universal properties of blowups and normalization [31, Proposition II.7.14, Exercise II.3.8], the map $v^{\prime}$ factors through $v:X^{+}\longrightarrow X$ (see Section missing 2). Let $g:X^{\prime}\longrightarrow X^{+}$ be the factoring map. We have the following isomorphism induced by $g^{*}\left(\mathcal{O}_{X^{+}}\right)=g^{-1}\mathcal{O}_{X^{+}}\otimes_{g^{-1}\mathcal{O}_{X^{+}}}\mathcal{O}_{X^{\prime}}\xrightarrow{\,\,\,\,\sim\,\,\,\,}\mathcal{O}_{X^{\prime}}$
@@
Indeed, a surjection of invertible sheaves is an isomorphism. Since $g$ is also projective [54, tag 01W7, Lemma 29.43.15] we have $g_{*}\left(\mathcal{O}_{X^{\prime}}\right)=\mathcal{O}_{X^{+}}$ by [31, proof of Corollary III.11.4] and [54, tag 01W7, Lemma 29.43.4]. Therefore, for every $n\in{\mathbb{Z}}_{>0}$ we have the following commutative diagram of natural maps.

### v3

@@ Abstract.
We provide explicit descriptions for the rational powers and Rees valuations of several classes of ideals invariant under natural actions of tori and products of general linear groups, in terms of polyhedra and lattice points. This allows us to show that a version of Mustaţă-Takagi’s summation formula for multiplier ideals also holds for the rational powers of these ideals. Moreover, for arbitrary ideals in normal domains that are finitely generated over algebraically closed fields, we prove a weaker version of this formula that holds for sufficiently large rational numbers.
@@ 1. Introduction
Rees valuations and rational powers of ideals play a central role in many results in algebraic geometry and commutative algebra [36, 39, 20, 13, 12, 46, 27, 43, 3, 26, 52]. These two sets of invariants are so closely related that each uniquely determines the other (see Section missing 2). However, explicitly computing them is a difficult task because it requires determining the order valuations for every prime divisor arising in the exceptional divisor of the normalized blowup of the ideal [50]. In fact, complete descriptions of Rees valuations in the literature are limited to monomial ideals, where they correspond to supporting hyperplanes of Newton polyhedra [38, Theorem 10.3.5]. One of the primary goals of this paper is to provide explicit descriptions of both Rees valuations and rational powers for several classes of ideals in polynomial rings that are invariant under natural actions of tori and products of general linear groups.
Originally considered by Samuel and Rees in the 1950s [54, 50], the notion of rational powers of an ideal $I$ in a Noetherian domain $R$ was introduced by Lejeune-Jalabert and Teissier [45, §4.2]. The power $\overline{I^{w}}$ for $w\in\mathbb{Q}_{\geqslant 0}$ is defined as $\overline{I^{w}}:=\cap_{v}\{x\mid v(x)\geqslant wv(I)\}$, where the intersection is taken over all the discrete valuations of $\operatorname{Quot}(R)$ that are nonnegative in $R$, or alternatively, over the Rees valuations of $I$. This construction provides a powerful tool for understanding the structure of filtrations of ideals such as integral closures [36, 39], symbolic powers [26], Frobenius powers [52], and multiplier ideals [44, §9.6]. Despite its historical roots, rational powers have seen limited development, as noted by Knutson in 2010, who observed that this “beautiful construction has sat essentially unused in the literature for over 50 years now" [42].
Let $X=X_{m\times n}$ be a generic $m\times n$ matrix, meaning its entries are distinct variables. The polynomial ring $\mathbb{C}[X]$ carries a natural action of $\mathrm{GL}_{m}(\mathbb{C})\times\mathrm{GL}_{n}(\mathbb{C})$ defined by $(A,B)\cdot X=A^{-1}XB$, where $A\in\mathrm{GL}_{m}(\mathbb{C})$ and $B\in\mathrm{GL}_{n}(\mathbb{C})$. The ideals in the polynomial ring $\mathbb{C}[X]$ that are invariant under this action have been extensively studied in the literature, see e.g. [17, 18, 6, 10, 35, 40, 21, 7]. The combination of representation theory and the theory of algebras with straightening laws has led to a wide range of explicit results about the structure and properties of these ideals. If one considers a symmetric generic matrix $Y=Y_{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{C})$-action on $\mathbb{C}[Y]$ given by $A\cdot Y=AYA^{t}$, or a skew-symmetric generic matrix $Z=Z_{n\times n}$ with $\operatorname{GL}_{n}(\mathbb{C})$-action on $\mathbb{C}[Z]$ given by $A\cdot Z=AZA^{t}$, one obtains analogous well-studied invariant ideals, see e.g. the references above and [2, 1, 19, 14].
We introduce the concept of Rees package, which is a combinatorial characteristic-free construction that encapsulates the description of both Rees valuations and rational powers in terms of polyhedral data (see Section missing 3). In our first theorem, we show that the classes of invariant ideals described above admit such a Rees package. Furthermore, in arbitrary characteristic, we show that monomial ideals in affine semigroup rings, sums of products of determinantal ideals of generic, symmetric, and skew-symmetric matrices, and products of determinantal ideals of Hankel matrices, also admit a Rees package.
@@
Let $\mathbb{k}$ be an arbitrary field. The following classes of ideals admit a Rees package. In particular, their rational powers and Rees valuations can be explicitly described in terms of lattice points and supporting hyperplanes of polyhedra.
@@
    (Theorem 3.2, §missing 3.2) Sums of products of determinantal ideals in $\mathbb{k}[X_{m\times n}]$ of a generic matrix $X_{m\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{C}[X_{m\times n}]$.
@@
    (Theorem 3.3, §missing 3.3) Sums of products of determinantal ideals in $\mathbb{k}[Y_{n\times n}]$ of a generic symmetric matrix $Y_{n\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{k}[Y_{n\times n}]$
@@
    (Theorem 3.4, §missing 3.4) Sums of products of ideals of Pfaffians in $\mathbb{k}[Z_{n\times n}]$ of a skew-symmetric generic matrix $Z_{n\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{C}[Z_{n\times n}]$.
@@
    (Theorem 3.5) Products of determinantal ideals in $\mathbb{k}[W]$ of a Hankel matrix of variables $W$.
In [49], Mustaţă showed a summation formula for multiplier ideals of pairs ideal sheaves, and this formula was later generalized by Takagi [56] and by Jow-Miller [41]. Mustaţă-Takagi summation formula raises the natural question of whether similar identities hold for symbolic powers and rational powers of ideals. The former was established in [32, 31] and the latter in [4] for monomial ideals. Thus we propose the following conjecture (cf. [4, Question 2.4]).
@@
We devote the last two sections of the paper to tackle this conjecture. First, we note that the restriction on the field is partially justified by a related question (see Section missing 4 and Section missing 4). We also note that the use of two different $\mathbb{k}$-algebras is essential in this conjecture, as the summation formula for rational powers does not hold in general if instead one considers two ideals in the same ring, see Section missing 4.
In Section missing 4 we show that if two ideals $I\subset R$ and $J\subset S$ admit a Rees package, e.g. the invariant ideals in Theorem A, then $IT+JT$ does as well. Moreover, a Rees package for $IT+JT$ can be obtained by ‘gluing’ those of $I$ and $J$. In particular, this shows that the Rees valuations of $IT+JT$ are obtained by pairing the ones of $I$ and $J$ (see Section missing 4). This result allows us to compute rational powers of more classes of ideals, see for instance Section missing 4.
In our next theorem, we use Section missing 4 to settle Section missing 1 for ideals with a Rees package without restrictions on the field.
Theorem B (Theorem 4.1, Section missing 4).
Adopt the notation in Section missing 1 with $\mathbb{k}$ an arbitrary field. The conclusion of Section missing 1 holds if $I$ and $J$ admit a Rees package. In particular, this holds if $I$ and $J$ belong to any of the classes of ideals in Theorem A, and not necessarily both from the same class.
In the following theorem, we provide further evidence for Section missing 1. Here we prove an inclusion for arbitrary ideals and $w\gg 0$. This result can be seen as a weaker version of Section missing 1 since we always have
see Section missing 5.
Theorem C (Theorem 5.1).
Following the notation in Section missing 1, there exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}_{\geqslant 0}$ then
@@ 2. Rational powers and Rees valuations
In this section we review and extend some relevant facts about rational powers and Rees valuations. Our main references in this subject is [38]. We note that several of the statements in this section have been previously established for the integral closures of powers of ideals. Our goal is to show that these properties continue to hold in the more general setting of rational powers and to provide references for the corresponding assertions. We restrict our exposition to the following setup where all of our results occur.
@@
If $w$ is an integer, the statement follows from [38, Proposition 10.5.5]. Now let $w$ be arbitrary and write $w=\frac{p}{q}$. If $x\in\overline{I^{w}}$, then $(W^{-1}x)^{q}=W^{-1}x^{q}\in W^{-1}\overline{I^{p}}=\overline{(W^{-1}I)^{p}}$ and so $W^{-1}x\in\overline{(W^{-1}I)^{w}}$. On the other hand, if $y\in\overline{(W^{-1}I)^{w}}$, then $y^{q}\in\overline{(W^{-1}I)^{p}}=W^{-1}\overline{I^{p}}$. Therefore there exists $a\in W$ such that $ay^{q}\in\overline{I^{p}}$. The latter implies $(ay)^{q}\in\overline{I^{p}}$ and so $ay\in\overline{I^{w}}$. Thus $y\in W^{-1}\overline{I^{w}}$, finishing the proof. ∎
A valuation is discrete if it is integer valued. Integral closures and rational powers can also be defined via discrete valuations: by [38, Theorem 6.8.3(3)] we have that $x\in\overline{I}$ if and only if $x\in IV$ for every discrete valuation ring $R\subseteq V\subseteq\operatorname{Quot}(R)$. Moreover, the latter condition can be verified by using finitely many valuations, the Rees valuations of $I$.
@@
Let ${\operatorname{\mathscr{RV}}}(I)=\{v_{1},\dots,v_{r}\}$ and set $e:={\operatorname{lcm}}(v_{1}(I),\dots,v_{r}(I))$ (or any multiple of it). For every $w\in{\mathbb{Q}}_{\geqslant 0}$ and $n:=\lceil we\rceil$ we have $\overline{I^{w}}=\overline{I^{\frac{n}{e}}}$ (see [38, Proposition 10.5.5]). Moreover, $\overline{I^{\frac{1}{e}}}=\sqrt{I}$ [38, Example 10.5.4].
@@
The second equality is proved in [38, Discussion 10.1.3] and the inclusion $\supseteq$ in the first one is trivial. It remains to show the first inclusion $\subseteq$. For each $V\in{\operatorname{\mathscr{RV}}}(I)$, let $\eta_{V}$ be its maximal ideal. From Section missing 2 we obtain $\overline{I^{w}}=\cap_{V\in{\operatorname{\mathscr{RV}}}(I)}(\eta_{V}^{\lceil wv(I)\rceil}\cap R)$. The result follows as each ideal $\eta_{V}^{\lceil wv(I)\rceil}\cap R$ is $\mathfrak{p}_{V}$-primary. ∎
@@
The implication (i) $\Rightarrow$ (ii) follows from Theorem 2.1, and (ii) $\Rightarrow$ (iii) from Section missing 2. It remains to show (iii) $\Rightarrow$ (i), but this is content of [38, Theorem 10.1.6], finishing the proof. ∎
@@
The equivalence (i) $\Leftrightarrow$ (iii) in Section missing 2 shows that if one knows the Rees valuations of an ideal, then one can compute its rational power; and conversely, if one knows the rational powers of the ideal, then one can determine its Rees valuations.
[…]
@@
For a particular example, consider $R$ a polynomial ring over a field and $H_{1},\ldots,H_{t}$ subspaces of $R_{1}$. For each $i$ let $I_{i}$ be the ideal generated by $H_{i}$ and set $I=I_{1}\cdots I_{t}$. The powers of $I$ are all integrally closed and irredundant primary decompositions of them having the same format above can be found in [16, Theorem 3.2]. Therefore, these primary decompositions can be used to explicitly describe the Rees valuations and rational powers of $I$.
@@
We begin with (1). By [38, Theorem 2.3.2], the algebra $C$ is ${\mathbb{Z}}_{\geqslant 0}$-graded and its $e$th-Veronese is the integral closure of $A_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{n}u^{en}$ in $B_{e}:=R[u^{e}]$, which is $C_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{n}}u^{en}$ by [38, Proposition 5.2.1]. Write $C=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}J_{n}u^{n}$ and fix $n$. If $x\in J_{n}$, then $x^{e}\in J_{en}=\overline{I^{n}}$, and so $x\in\overline{I^{\frac{n}{e}}}$. On the other hand, if $x\in\overline{I^{\frac{n}{e}}}$, then $x^{e}\in\overline{I^{n}}$, and so $(xu^{n})^{e}\in C$. Thus, $xu^{n}$ is integral over $A$ and then $xu^{n}\in C$. It follows that $x\in J_{n}$.
We continue with (2). Note that $A=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{\lceil\frac{n}{e}\rceil}u^{n}$ and is Noetherian. By Rees’ theorem [51] (see also [38, Theorem 9.1.2]), there exists $k\in{\mathbb{Z}}_{>0}$ such that, for every integer $n\geqslant ek$, we have
@@
Finally we prove (3). By (2) and [28, Lemma 13.10], there exists $m\in{\mathbb{Z}}_{>0}$ such that $\overline{I^{mt}}=\left(\overline{I^{m}}\right)^{t}$ for every $t\in{\mathbb{Z}}_{\geqslant 0}$. Set $C_{m}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{mt}}u^{emt}$. From (2), it follows that $C$ is a finitely generated $C_{m}$-algebra, and since $C_{m}\subset C$ is an integral extension, by [38, Lemma 2.1.9], $C$ is a finitely generated $C_{m}$-module. The statement now follows by considering $r\in{\mathbb{Z}}_{>0}$ to be the largest degree of an element in a finite set of generators of $C$ as a $C_{m}$-module. ∎
We continue with the following statement that allows for base change in rational powers (cf. [38, Corollary 19.5.2]). A ring homomorphism $f:R\to S$ is normal if it is flat and if for every $\mathfrak{p}\in{\operatorname{Spec\,}}(R)$ and field extension $\mathbb{K}$ of $R_{\mathfrak{p}}/\mathfrak{p}R_{\mathfrak{p}}$ the ring $\mathbb{K}\otimes_{R}S$ is normal.
@@
Let $e_{1}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(I)\}\right)$, $e_{2}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(IS)\}\right)$, and $e:={\operatorname{lcm}}(e_{1},e_{2})$. Let $u,A,B,C$ be as in Section missing 2 defined using the latter $e$. By [38, Theorem 19.5.1], $C\otimes_{R}S$ is the integral closure of $A\otimes_{R}S$ in $B\otimes_{R}S$. Under the isomorphism $B\otimes_{R}S\cong S[u]$, $A\otimes_{R}S$ corresponds to $S[(IS)u,(IS)u^{2},\ldots,(IS)u^{e}]$ and $C\otimes_{R}S$ to $\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}(\overline{I^{\frac{n}{e}}}S)u^{n}$. The result follows by applying Section missing 2(1) to $IS$. ∎
@@
In this subsection we introduce the notion of rational powers of ideal sheaves in normal varieties over algebraically closed fields, inspired by [45, Définition 4.2.2] and [44, Definition 9.6.2].
@@
Let $X$ be a normal variety (integral, separated, and of finite type) over an algebraically closed field $\mathbb{k}$ and let $\mathcal{I}$ be a nonzero $\mathcal{O}_{X}$-ideal.
@@
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor. Note that $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. The integral closure of $\mathcal{I}$, denoted by $\overline{\mathcal{I}}$, is the $\mathcal{O}_{X}$-ideal ¯$\overline{\mathcal{I}}:=v_{*}\left(\mathcal{O}_{X^{+}}(-E)\right)$ (cf. [44, Definition 9.6.2]). Moreover, for $w\in\mathbb{Q}_{\geqslant 0}$, we call the $\mathcal{O}_{X}$-ideal
@@
If $X$ in Section missing 2 is affine, so that $\mathcal{I}=\tilde{I}$ for some $\mathbb{k}[X]$-ideal $I$, then from Section missing 2 it follows that $\overline{I^{w}}=\Gamma(X,\overline{\mathcal{I}^{w}})$ are the rational powers of $I$ as in Section missing 2.
@@
We have the following alternative way of computing rational powers similar to [44, Remark 9.6.4] and [45, Théorème 2.1]. We use the definition of projective morphism from [25], which is more general than the one of [34] (see [55, tag 01W7, Lemma 29.43.3]), and includes blowups of arbitrary varieties along ideal sheaves [55, tag 01OF, Lemma 31.32.13].
@@
By the universal properties of blowups and normalization [34, Proposition II.7.14, Exercise II.3.8], the map $v^{\prime}$ factors through the normalized blowup $v:X^{+}\longrightarrow X$. Let $g:X^{\prime}\longrightarrow X^{+}$ be the factoring map. Let $E$ be such that $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. We have the following isomorphism induced by $g^{*}\left(\mathcal{O}_{X^{+}}\right)=g^{-1}\mathcal{O}_{X^{+}}\otimes_{g^{-1},\mathcal{O}_{X^{+}}}\mathcal{O}_{X^{\prime}}\xrightarrow{\,\,\,\,\sim\,\,\,\,}\mathcal{O}_{X^{\prime}}$
@@
Indeed, a surjection of invertible sheaves is an isomorphism. Since $g$ is also projective [55, tag 01W7, Lemma 29.43.15] we have $g_{*}\left(\mathcal{O}_{X^{\prime}}\right)=\mathcal{O}_{X^{+}}$ by [34, proof of Corollary III.11.4] and [55, tag 01W7, Lemma 29.43.4]. Therefore, for every $n\in{\mathbb{Z}}_{>0}$ we have the following commutative diagram of natural maps.